# Multidimensional Rubric-oriented Reward Model Learning via Geometric Projection Reference Constraints

### v2

@@ 2.1 Paradigm Evolution of Medical LLM Evaluation and Alignment
Medical LLM evaluation has evolved from "knowledge assessment" to "capability evaluation" and "scenario adaptation," but core limitations persist. Early methods (e.g. MedQA) (roy2024beyond) focus on medical factual knowledge via multiple-choice questions. They are easy to quantify but fail to evaluate medical-critical capabilities like reasoning, communication, and compliance, leading to a disconnect from real needs. Subsequent benchmarks (HealthBench) (arora2025healthbench) introduce scenario-based tasks with refined evaluation rules. However, they lack a "evaluation-training" closed-loop, limiting generalization across scenarios and dimensions. Reinforcement Learning with Verifiable Rewards (RLVR) shows potential in medical alignment, while extending it to real-world tasks is challenging, as evaluation relies on nuanced multi-criteria judgments rather than binary correctness. Recent structured rubric systems address multi-dimensional evaluation by rule-based or LLM-based rubric -level on-training scoring, which is either static or high-cost: (1) Rubicon-preview (huang2025reinforcement): A rubric library with tens of thousands of standards, extending verifiable reward RL to open domains. Limitations include general-task-oriented dimensions, poor rule reusability, and reliance on pre-defined rules (inconsistent evaluation). (2) RaR (gunjal2025rubrics) implements evaluation using instance-specific rubrics generated by large language models (LLMs), transforming rubrics from mere evaluation tools into reward mechanisms and bridging the gap between verifiable rewards and preference ranking. This LLM-as-Judge approach is also adopted in (ye2025self; xie2025auto), with additional rubrics learning processes, etc. However, they have several core limitations: first, rubric generation depends on real-time LLM-based scoring, which is costly and lacks sufficient compliance depth for medical scenarios; second, rubric standards are tightly coupled with specific tasks, necessitating re-generation when new evaluation dimensions are added; third, it over-relies on LLM-as-Judge during reinforcement learning training, leading to extremely high training costs and imposing strict requirements on both the consistency of the scoring LLM and the quality of training data standards.
@@ 3.4.2 Model Architecture and Training Process
  • Question-answer vectors: $\vec{v}^{i}_{qab}=\phi(q\oplus ab_{i,q})\in\mathbb{R}^{d}$, $\vec{v}^{i}_{qar}=\phi(q\oplus ar_{i,q})\in\mathbb{R}^{d}$, $\vec{v}^{i}_{qaw}=\phi(q\oplus aw_{i,q})\in\mathbb{R}^{d}$, where $\oplus$ denotes concatenation. Subscript $b$, $r$ and $w$ represent better, reference and worse sample, respectively.
  • Dimension description vector: $\vec{v}^{i}_{\text{Desc}}=\phi(\text{Desc}_{i})\in\mathbb{R}^{d}$.
(2) Dimension-Specific Projection Scoring
Answer vectors are projected onto $\vec{v}^{i}_{\text{Desc}}$ to calculate $L_{i}$ scores with dot product normalization to eliminate vector length bias:
  • Good sample score: $s^{i}_{ab}=\frac{\vec{v}^{i}_{qab}\cdot\vec{v}^{i}_{\text{Desc}}}{\|\vec{v}^{i}_{\text{Desc}}\|}$.
  • Medium sample score: $s^{i}_{ar}=\frac{\vec{v}^{i}_{qar}\cdot\vec{v}^{i}_{\text{Desc}}}{\|\vec{v}^{i}_{\text{Desc}}\|}$.
  • Poor sample score: $s^{i}_{aw}=\frac{\vec{v}^{i}_{qaw}\cdot\vec{v}^{i}_{\text{Desc}}}{\|\vec{v}^{i}_{\text{Desc}}\|}$.
(3) Pairwise Gap Modeling Loss
To capture the relative quality gaps between different model outputs, we employ the Bradley-Terry (BT) loss, a standard approach for pairwise comparison. The BT loss models the probability of item $i$ defeating item $j$ as $P(i>j)=\sigma(s_{i}-s_{j})$, where $s$ represents the quality score and $\sigma$ is the sigmoid function. Let the score difference be $\delta_{ij}=s_{i}-s_{j}$. We unify the loss as a Binary Cross-Entropy (BCE) format:
$f_{BT}(\delta_{ij})=-y\cdot\log(\sigma(\delta_{ij}))-(1-y)\cdot\log(1-\sigma(\delta_{ij}))$
where $y=1$ indicates the true outcome is $i$ defeats $j$, and $y=0$ otherwise. In our specific application, we are interested in the quality gaps for three predefined pairs: ($aw,ab$), ($ar,ab$), and ($aw,ar$). We define our quality scores $\theta$ as the model’s predicted scores, $s^{i}_{aw},s^{i}_{ab},s^{i}_{ar}$. For each pair, we model the probability of the first item being better than the second (i.e., $y=1$). Let the score difference be $\delta^{i}_{bw}=s^{i}_{ab}-s^{i}_{aw},\delta^{i}_{br}=s^{i}_{ab}-s^{i}_{ar},\delta^{i}_{rw}=s^{i}_{ar}-s^{i}_{aw}$. The loss for each pair is thus calculated as: $L_{BT_{bw}}^{i}=f_{BT}({\delta^{i}_{bw}})$, $L_{BT_{br}}^{i}=f_{BT}({\delta^{i}_{br}})$ and $L_{BT_{rw}}^{i}=f_{BT}({\delta^{i}_{rw}})$.
(4) Geometric Constraint Regularization
Based on medical cognitive logic, we hope that the projection score for medium sample lays between that of poor and good. The Geometric Constraint Regularization loss is introduced:
$L^{i}_{\text{GC}}=\left[L_{BT_{bw}}^{i}-(L_{BT_{rw}}^{i}+L_{BT_{br}}^{i})\right]^{2}$
This restricts extreme sample gap over-amplification, ensuring scores align with medical rules.
(5) Total Loss and Training Optimization
Total loss function for a query $q$ for training reward model is:
$\mathcal{L}_{\text{RM}}=\sum^{L}_{i=1}(L_{BT_{bw}}^{i}+L_{BT_{br}}^{i}+L_{BT_{rw}}^{i}+\lambda\cdot L^{i}_{\text{GC}})$
where $\lambda$ is the regularization coefficient.

### v3

@@ 2.1 Paradigm Evolution of Medical LLM Evaluation and Alignment
Medical LLM evaluation has evolved from "knowledge assessment" to "capability evaluation" and "scenario adaptation," but core limitations persist. Early methods (e.g. MedQA) (roy2024beyond) focus on medical factual knowledge via multiple-choice questions. They are easy to quantify but fail to evaluate medical-critical capabilities like reasoning, communication, and compliance, leading to a disconnect from real needs. Subsequent benchmarks (HealthBench) (arora2025healthbench) introduce scenario-based tasks with refined evaluation rules. However, they lack a "evaluation-training" closed-loop, limiting generalization across scenarios and dimensions. Reinforcement Learning with Verifiable Rewards (RLVR) shows potential in medical alignment, while extending it to real-world tasks is challenging, as evaluation relies on nuanced multi-criteria judgments rather than binary correctness. Recent structured rubric systems address multi-dimensional evaluation by rule-based or LLM-based rubric-level on-training scoring, which is either static or high-cost: (1) Rubicon-preview (huang2025reinforcement): A rubric library with tens of thousands of standards, extending verifiable reward RL to open domains. Limitations include general-task-oriented dimensions, poor rule reusability, and reliance on pre-defined rules (inconsistent evaluation). (2) RaR (gunjal2025rubrics) implements evaluation using instance-specific rubrics generated by LLMs, transforming rubrics from mere evaluation tools into reward mechanisms and bridging the gap between verifiable rewards and preference ranking. This LLM-as-Judge approach is also adopted in (ye2025self; xie2025auto; jacob2025qa), with additional rubrics learning processes, etc. However, they have several core limitations: first, rubric generation depends on real-time LLM-based scoring, which is costly and lacks sufficient compliance depth for medical scenarios; second, rubric standards are tightly coupled with specific tasks, necessitating re-generation when new evaluation dimensions are added; third, it over-relies on LLM-as-Judge during reinforcement learning training, leading to extremely high training costs and imposing strict requirements on both the consistency of the scoring LLM and the quality of training data standards.
@@ 3.4.2 Model Architecture and Training Process
  • Question–answer vectors: $\vec{ob}=\phi(q\oplus ab_{i,q})\in\mathbb{R}^{d}$, $\vec{or}=\phi(q\oplus ar_{i,q})\in\mathbb{R}^{d}$, $\vec{ow}=\phi(q\oplus aw_{i,q})\in\mathbb{R}^{d}$, where $\oplus$ denotes concatenation. Subscript $\vec{ob}$, $\vec{or}$ and $\vec{ow}$ represent better, reference and worse samples, respectively.
  • Prompt-anchor vector: $\vec{op}^{\,i}=\phi(q\oplus\text{Desc}_{i})\in\mathbb{R}^{d}$, used as the task-dimension axis for projections and ratio consistency in dimension $i$.
We use $\vec{op}^{\,i},\vec{ow},\vec{or},\vec{ob}$ as the embeddings for axis-aligned and perpendicular computations relative to the prompt-anchored task-dimension axis.
(2) Prompt-anchored Polar Decomposition
We construct a polar coordinate system with the prompt as axis. Let $\vec{op}^{\,i},\vec{ow},\vec{or},\vec{ob}$ denote last-token embeddings of prompt (dimension-$i$ axis), worse, reference and better. Define the axis unit $\hat{e}_{p}^{\,i}=\dfrac{\vec{op}^{\,i}}{\|\vec{op}^{\,i}\|+\varepsilon}$, where $\varepsilon>0$. For any sample vector $\vec{o}\in\mathbb{R}^{d}$, we decompose it into the axis-aligned scalar projection and the perpendicular component. This polar view makes the training objective geometrically complete: the axis-aligned term $o^{\parallel,i}$ encodes alignment along the task dimension (ensuring $\texttt{worse}<\texttt{reference}<\texttt{better}$ on the prompt axis), while the perpendicular term $\|\vec{o}^{\perp,i}\|$ encodes off-axis deviation (ensuring better/reference are less off-topic than worse). Optimizing both eliminates the unrecognized degree of freedom that appears when only one term is used.
$o^{\parallel,i}$ $=\vec{o}\cdot\hat{e}_{p}^{\,i},$
$\vec{o}^{\perp,i}$ $=\vec{o}-o^{\parallel,i}\,\hat{e}_{p}^{\,i}$
Here $\hat{e}_{p}^{\,i}$ is the unit prompt axis for dimension $i$; $o^{\parallel,i}$ is the scalar projection of $\vec{o}$ onto the axis; $\vec{o}^{\perp,i}$ is the component orthogonal to the axis; and axis deviation is measured by $\|\vec{o}^{\perp,i}\|$. This yields axis similarity via $o^{\parallel,i}$ and axis deviation via $\|\vec{o}^{\perp,i}\|$, jointly determining each sample’s polar placement relative to the prompt.
(3) Axis Similarity Ranking
We encode the semantic ordering $\texttt{worse}<\texttt{reference}<\texttt{better}$ along the prompt axis using pairwise logistic (Bradley–Terry style) ranking on scalar projections:
$\delta^{\parallel,i}_{br}=\vec{ob}\cdot\hat{e}_{p}^{\,i}-\vec{or}\cdot\hat{e}_{p}^{\,i},\quad\delta^{\parallel,i}_{rw}=\vec{or}\cdot\hat{e}_{p}^{\,i}-\vec{ow}\cdot\hat{e}_{p}^{\,i},$
Here $\delta^{\parallel,i}_{br}$ and $\delta^{\parallel,i}_{rw}$ denote axis-margin differences between the pairs (better, reference) and (reference, worse) in dimension $i$.
$L_{\parallel}^{i}=-\log\sigma\big(\delta^{\parallel,i}_{br}\big)\;-\;\log\sigma\big(\delta^{\parallel,i}_{rw}\big).$
Here $\sigma(\cdot)$ is the logistic function and $L_{\parallel}^{i}$ is the axis-similarity ranking loss for dimension $i$.
(4) Perpendicular Deviation Ranking
We simultaneously enforce decreasing deviation from the axis: $\texttt{worse}>\texttt{reference}>\texttt{better}$ in $\|\vec{o}^{\perp,i}\|$. Define
$\Delta^{\perp,i}_{wr}=\|\vec{ow}^{\perp,i}\|-\|\vec{or}^{\perp,i}\|,\quad\Delta^{\perp,i}_{rb}=\|\vec{or}^{\perp,i}\|-\|\vec{ob}^{\perp,i}\|,$
Here $\Delta^{\perp,i}_{wr}$ and $\Delta^{\perp,i}_{rb}$ denote the margins in perpendicular deviation between (worse, reference) and (reference, better) in dimension $i$.
$L_{\perp}^{i}=-\log\sigma\big(\Delta^{\perp,i}_{wr}\big)\;-\;\log\sigma\big(\Delta^{\perp,i}_{rb}\big).$
Here $L_{\perp}^{i}$ is the perpendicular-deviation ranking loss for dimension $i$. This ranking encourages reference/better to align closer to the prompt axis than worse.
(5) Total Objective
The final objective balances axis similarity and deviation control:
$\mathcal{L}_{\text{RM}}^{i}=\alpha\,L_{\parallel}^{i}+(1-\alpha)\,L_{\perp}^{i},\quad\alpha\in(0,1).$
Here $\mathcal{L}_{\text{RM}}$ is the total objective and $\alpha$ balances axis similarity and deviation control. The prompt-anchored polar view separates “along-axis semantics” from “off-axis deviation”, yielding interpretable gradients and robust geometry. Stability is ensured by saturating $\log\sigma(\cdot)$ on large margins and unit-axis normalization and $\varepsilon$ safeguards in norms.